OpenStudy (anonymous):

Evaluate. 7!/4!3! 1 35 210

Parth (parthkohli):

\[\Large \color{purple}{\rightarrow {7 * 6 * 5 * 4 * 3 * 2 * 1 \over 4 * 3^2 * 2^2 * 1 } }\]

OpenStudy (anonymous):

no need to expand... just use 7! = 7*6*5*4! and cancel out the 4!

Parth (parthkohli):

Cancel the terms that exist in both numerator and denominator.

Parth (parthkohli):

I was explaining it in the most basic way.

OpenStudy (callisto):

\[\frac{7!}{4!3!} = \frac{7\times 6\times 5\times 4!}{4!3!}=\frac{7\times 6\times 5}{3 \times 2 \times 1}\]\[=7 \times 5 =?\]

OpenStudy (anonymous):

35

OpenStudy (callisto):

Yup :)